A piece of pipe is 50 inches long and cut into 3 pieces. The longest piece is 10 inches longer than the middle piece,while the shortest piece is 5 inches less than the middle piece. Please help me find the equation. I have the solution: 25 inches, 15 inches and 10 inches, but I cannot figure out the equation used to get these answers.
Thanks,

Answer by mananth(16946)

let the middle piece be x cm
the flongest piece will be x+10
the shortest piece = x-5
The sum of the pieces = 50, the total length
..
x+x+10+x-5=50
3x+5=50
add -5 to both sides
3x-5+5 = 50-5
3x = 45
divide by 3
3x/3 = 45/3
x=15 inches the middle piece
largest will be x+6 = 15+6 =21 inches
the shortest will be x-5 = 15-5 = 10 inches